# Autonomous Vehicles Using Multi-Agent Reinforcement Learning for Routing Decisions Can Harm Urban Traffic

### v1

@@ Appendix C SUMO
In this study, we consider that SUMO operates under deterministic conditions. Specifically, if all vehicles select the same routes across consecutive iterations, their travel times will remain identical to those observed in the previous iterations. The videos from where the screenshots were taken are included in the supplementary material.
(a) System optimal user equilibrium.
(b) Suboptimal user equilibrium.
Figure 8: SUMO screenshots illustrate the network under the two distinct equilibria. Red vehicles represent human drivers and yellow ones represent AVs.
Appendix D Cumulative time difference
@@
Positive cumulative travel time differences, i.e. $c_{t}>0$, mean that on average, human agents after mutation experience longer travel times, thus the impact of AVs introduction on them is negative. This occurred in the system optimal user equilibrium scenario for each algorithm tested. In the suboptimal user equilibrium, $c_{t}<0$ for each algorithm. This means that human agents decrease their travel times compared to the ones they were experiencing before the AVs introduction. This can be attributed to the fact that humans follow the route that has priority (route 1) and these values are negligible.
@@
where lh stands for the last episode before introducing AVs. For each algorithm and each equilibrium, there are 10 experiment replications, set $\mathcal{H}$ contains 12 human agents, mutation started in the 201-st episode and experiments took place until the 8300 episode. The values in the Table 3 are cumulative travel time differences and the standard deviations over experiments.
Appendix E Travel times
Figure 9: Mean travel times of AVs, human agents, and both in the simulation. The two plots on the left show that IPPO and MAPPO algorithms achieve shorter travel times for both AVs and human agents, as they converge to more optimal solutions, according to Figure 4. In the two plots on the right, that represent the scenario where humans adapt their options (Section 3), AVs achieve smaller travel times using the same algorithms. Human drivers gain no benefit. The error bars represent the results of three distinct experiments conducted for each algorithm in each equilibrium.
Appendix F Adaptation
Figure 10: 8000 episodes (days) of training for selected MARL algorithms while humans adapt their choices. The trivial solution is found only by IPPO and MAPPO in the suboptimal user equilibrium. Other algorithms remain probabilistic after and fail to converge to optimal policy. Suboptimal equilibrium is reproduced after 1000 days.
Appendix G Results for smaller AV fleets
@@
Appendix H Mean rewards
@@
Figure 13: Mean rewards of AV agents, human agents, and both combined in the suboptimal user equilibrium. IPPO and MAPPO achieve the lowest average rewards in the testing phase and during the training their reward is less oscillatory than the other algorithms. The rewards in this scenario are higher than those in Figure 12, as the system is under the suboptimal use equilibrium.
Appendix I Action shifts

### v2

@@ Appendix A Experiments
Hardware.
Our experiments were carried out on our institution’s computing nodes with resources allocated as listed in Table 4.
Table 4: Summary of the hardware used for the experiments.
Component Specification
CPU Intel(R) Xeon(R) Gold 5122 CPU, 3.60GHz
GPU NVIDIA GeForce RTX 2080
RAM 40 GB allocated per job
Operating system Ubuntu 24.04.1 LTS
Job scheduler SLURM
SUMO version 1.18.0
Execution time. Experiment compute time depends on the MARL algorithm. We share the computation time of some representative cases in Table 5.
Table 5: Computation time of representative experiments.
Traffic network Algorithm Runtime (min)
TRY IQL $\sim$64
TRY QMIX $\sim$80
TRY VDN $\sim$80
TRY IPPO $\sim$200
TRY MAPPO $\sim$200
TRY ISAC $\sim$220
TRY MASAC $\sim$220
Ingolstadt IPPO $\sim$7200 (120 hrs)
Overall, the experiments described in the paper required approximately 2,000 hours of computation. Full research required up to twice as much computation in total, due to preliminary testing, errors, and our curiosity-driven exploration.
Appendix B Plots
In all plots, except Figure 5, the lines represent the mean across multiple replications for each episode, and the error bars indicate the 15th and 85th percentile values. In Figure 3, the line depicts the mean across replications, while the error band accounts for variations due to different agent start times and the use of various algorithms, as described in Section 5.1. All plotted data smoothed using a moving average method with a window size of value 10.
@@ Appendix C SUMO
SUMO is the traffic simulator used in this study and is licensed under the Eclipse Public License Version 2.0 (EPL V2). Throughout our experiments, SUMO operates under deterministic conditions, except Figure 4c, where we explicitly introduce non-determinism. Under deterministic settings, if all vehicles select the same routes across consecutive iterations, their travel times remain identical across runs.
In the SUMO screenshots below, red vehicles represent human drivers, and yellow vehicles depict AVs. The SUMO network file used to simulate the TRY network is provided in the supplementary material, along with simulation videos.
(a) System optimal user equilibrium (see Figure 1)a.
(b) Suboptimal user equilibrium (see Figure 1)b.
Figure 6: SUMO screenshots illustrate the network under the two distinct equilibria.
Figure 7: Sumo screenshot from the Ingolstadt network.
Appendix D Ingolstadt network experiment
The Ingolstadt network consists of 850 links, with demand data sourced from InTAS [34] and used in the RESCO benchmark [39]. This demand includes 1,000 agents. In our experiment, we consider that 400 human agents transitioned to using AVs in episode 500. Each agent learns to select the optimal path from four available alternatives. These alternative paths are generated using Janux software [1], which is integrated into the RouteRL framework.
As shown in Figure 5, initially, when there are only humans, the system remains stable, exhibiting constant average travel time. However, starting at episode 500, the introduction of AVs destabilizes the system, leading to oscillatory and increased patterns in travel times for human agents.
Appendix E Brief theoretical analysis of MARL convergence
This section outlines the theoretical barriers to the convergence of the MARL algorithms employed in this paper. Although empirical performance is often emphasized in MARL, for example in [17, 42], theoretical convergence guarantees are often lacking. Moreover, in recent benchmarks, agents are trained for millions of timesteps [17, 47].
In IL methods, such as IQL, each agent treats others as part of the environment. This breaks the Markov assumption and leads to non-stationarity, as the transition dynamics $P(s^{\prime}|s,a_{i})$ are no longer fixed, due to the evolving policies of other agents. As a result, even though IQL is a baseline used in MARL problems that often works well in practice, it lacks theoretical convergence guarantees [47, 19, 36].
Value factorization algorithms, such as VDN and QMIX assume that the joint action-value function $Q_{tot}(s,a)$ can be decomposed into individual functions $Q_{i}(s,a_{i})$, either additively (VDN) or under monotonicity constraint (QMIX). This assumption is violated in many coordination tasks, especially those evolving non-monotonic utility landscapes [41]. However, this property does not guarantee non-convergence.
As far as the SAC algorithm is concerned prior work has mentioned that MASAC cannot scale easily to the large joint action spaces inherent in multi-agent settings [67]. However, to the best of our knowledge, there is no proof to support this claim.
Appendix F Cumulative time difference
@@
Positive cumulative travel time differences, i.e. $c_{t}>0$, mean that on average, human agents after mutation experience longer travel times, thus the impact of AVs introduction on them is negative. This occurred in the system’s optimal user equilibrium scenario for each algorithm tested. In the suboptimal user equilibrium, $c_{t}<0$ for each algorithm. This means that human agents decrease their travel times compared to the ones they were experiencing before the introduction of AVs. This can be attributed to the fact that humans follow the route that has priority (route 1) and these values are negligible.
@@
where lh stands for the last episode before introducing AVs. For each algorithm and each equilibrium, we conducted 10 experiment replications. The set $\mathcal{H}$ contains 12 human agents. Mutation began in the 201-st episode and experiments continued until episode 8300. The values in the 6 are cumulative travel time differences and the standard deviations over experiments.
Appendix G Critical fleet analysis
Figure 8: With an increasing number of agents, a fraction of those, who learn the optimal policy decreases, achieving similar efficiency to random choice (equilibrium 0: system optimal - user equilibrium - Figure 1a, equilibrium 1: suboptimal user equilibrium - Figure 1b). This appears to happen with as few as four agents for some of the tested algorithms (ISAC, IQL) that failed to converge in Figure 4. See Appendix A for the experiments’ setup.
Appendix H Centralization and adaptation
Figure 9: Top. IPPO and MAPPO still converge if the adaptation term is added to the model (dotted lines). However, the system diverts from the optimal state to the suboptimal, where the total costs are higher. Bottom. The centralized version of IPPO can, in some cases, accelerate convergence reducing the time required to reach an optimal solution, see Appendix I for exceptions. However, this reduction is limited to a few days. See A for the experiments’ setup.
Appendix I Limits of the centralized case
Figure 10: In the system optimal user equilibrium case (1a), the centralized IPPO converges to the suboptimal solution, showing the limitations of the centralized case. This contrasts with the result shown in Figure 9 bottom, where the centralized case converges to the optimal solution faster than the decentralized version of the algorithm.
Appendix J Travel times
Figure 11: Mean travel times of AVs, human agents, and both in the simulation. The two plots on the left show that IPPO and MAPPO algorithms achieve shorter travel times for both AVs and human agents, as they converge to more optimal solutions, according to Figure 4(a, b). In the two plots on the right, that represent the scenario where humans adapt their options (2), AVs achieve smaller travel times using the same algorithms. Human drivers gain no benefit. The error bars represent the standard deviation between three distinct experiments conducted for each algorithm in each equilibrium.
Appendix K Adaptation
Figure 12: 8000 episodes (days) of training for selected MARL algorithms while humans adapt their choices. The trivial solution is found only by IPPO and MAPPO in the suboptimal user equilibrium. Other algorithms remain probabilistic after and fail to converge to optimal policy. The optimal solution in the suboptimal equilibrium is achieved after 1000 days. The error bars depict the 15th and 85th percentile values across replications and the solid line indicates the mean value across replications.
Appendix L Results for smaller AV fleets
@@ Appendix I Limits of the centralized case
Appendix M Mean rewards
@@
Figure 15: Mean rewards of AV agents, human agents, and both combined in the suboptimal user equilibrium. IPPO and MAPPO achieve the lowest average rewards in the testing phase and during the training their reward is less oscillatory than the other algorithms. The rewards in this scenario are higher than those in 14, as the system is under the suboptimal use equilibrium.
Appendix N Action shifts